OpenStudy (anonymous):

For 2x – 3y = –10, when x = 4, What is the value of y?

OpenStudy (anonymous):

negative two-thirds two-thirds 6 –6

Parth (parthkohli):

When you know what \(x\) is and you have to solve for \(y\), you should get a simple equation.

Parth (parthkohli):

\[2(4) - 3y = -10 \]Start by plugging 4 into \(x\).

Parth (parthkohli):

Now... what is \(2(4)\)?

OpenStudy (anonymous):

6-3y=-10

OpenStudy (anonymous):

be rigth back

Parth (parthkohli):

Remember:\[ 2(4) = 2\times 4 \ne 2 + 4\]

Parth (parthkohli):

2(4) is 2 x 4, but it is not 2 + 4.

OpenStudy (anonymous):

:)